A family flies from Kelowna to Chicago. The first flight takes them from Kelowna to Toronto, a distance of 3,060 km. The flight took 4 hours and 27 minutes. The family then had a layover at the Toronto airport for 2 hours and 12 minutes. Their next flight took them to Chicago, 700 km away, at a speed of 790 km/h. What is the average speed of their entire travel?
1
Expert's answer
2020-11-30T14:54:37-0500
"V=\frac{S}{T}\\V=\frac{3060+700}{\frac{4(60)+27}{60}+\frac{2(60)+12}{60}+\frac{700}{790}}\\V=499\frac{km}{h}"